# Combining Evidence Across Filtrations

### v4

@@ 1.3 Related work on combining e-values (and p-values)
Many recent papers highlight how e-values, unlike p-values, can be combined seamlessly, even when they are arbitrarily dependent (e.g., Vovk and Wang,, 2021; Wang and Ramdas,, 2022; Grünwald et al.,, 2024). In particular, Vovk and Wang, (2021) show that the arithmetic mean function, which straightforwardly yields a valid combined e-value, dominates any other symmetric “e-merging” function. This is useful when combining two or more different e-values for a hypothesis $\mathcal{P}$ and each e-value is constructed using the same data sample $X_{1},\dotsc,X_{n}$. In the sequential context, given that an e-process in a filtration $\mathbb{F}$ is an e-value at any $\mathbb{F}$-stopping time, any two e-processes in the same filtration $\mathbb{F}$ for a (shared) null hypothesis $\mathcal{P}$ can also be combined seamlessly by averaging. The main challenge we highlight in this work is when two e-processes are defined in two different filtrations, such that one (or both) e-processes must first be lifted into a common filtration before the averaging step.
Much of the literature also concerns cases where the e-values exhibit various dependence structures that can be exploited. Generally, taking the product of e-values is a (weakly) dominating e-merging strategy (Vovk and Wang,, 2021, proposition 4.2) when the e-values either are independent or form a supermartingale. Taking the product is a useful strategy in sequential settings involving continuous monitoring (Howard et al.,, 2021; Waudby-Smith and Ramdas,, 2024) and in meta-analysis (ter Schure and Grünwald,, 2022; Grünwald et al.,, 2024). In this work, we focus on cases where the two e-processes are constructed using the same data sequence, so the product strategy does not apply. Others study e-values for multiple hypotheses with different, including arbitrary, dependence structures (e.g., Xu et al.,, 2021; Wang and Ramdas,, 2022; Ignatiadis et al.,, 2024). We do not generally consider such multiple testing scenarios.
Finally, there are various strategies for combining p-values (see, e.g., Vovk and Wang,, 2020). Yet, Vovk et al., 2022b ’s recent results suggest that any admissible combination function for p-values must go through e-values. Thus, we primarily focus on e-processes, which are also the fundamental notion of evidence in SAVI (Ramdas et al.,, 2020, 2023). That said, we will additionally discuss the combination of e-processes with p-processes, which are the anytime-valid counterpart to p-values, as a preliminary step to our main result; this task is related to Ignatiadis et al., (2024)’s study on combining (non-sequential) e-values and p-values for multiple testing. As we shall see later, p-processes can be lifted “freely” across filtrations, and they appear in an intermediary step when combining e-processes across filtrations. The underlying connection is made via p-to-e calibrators, which we review in Section 2.2.
@@ 2.1 SAVI part 2: P-processes, sequential tests, and Ville’s inequality
Continuing from our review of SAVI in Section 1.1, we now define p-processes, which are the sequential generalization of p-values. Given a family of distributions $\mathcal{P}$ and a filtration $\mathbb{F}$, we say that a $[0,1]$-valued $\mathbb{F}$-process $(\mathfrak{p}_{t})_{t\geq 0}$ is a p-process for $\mathcal{P}$ in $\mathbb{F}$ if, for any $\alpha\in(0,1)$,
@@
A p-process readily yields the level-$\alpha$ sequential test $\phi_{t}=\mathds{1}\left(\mathfrak{p}_{t}\leq\alpha\right)$ for each $\alpha\in(0,1)$. An e-process also yields a sequential test via Ville’s inequality (Ville,, 1939), which was originally shown for test supermartingales but has since been extended to e-processes (Ramdas et al.,, 2020). The inequality states that, for any e-process $\mathfrak{e}=(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in $\mathbb{F}$, the probability under $\mathcal{P}$ that $\mathfrak{e}$ exceeds a large threshold is small at any $\mathbb{F}$-stopping time: for any $\alpha\in(0,1)$,
@@
the same filtration (Vovk and Wang,, 2020)
@@
Table 1: A comparison of evidence measures in SAVI: e-processes and p-processes. Both the e-process $(\mathfrak{e}_{t})_{t\geq 0}$ and the p-process $(\mathfrak{p}_{t})_{t\geq 0}$ are adapted to some underlying filtration $\mathbb{F}$, in which a stopping time $\tau$ can be defined. In the first row, $P\in\mathcal{H}_{0}$ denotes any member of the composite null hypothesis $\mathcal{H}_{0}$. The last row summarizes the main takeaways of Section 3.
@@ 2.2 Calibrating e-processes and p-processes
Combining evidence in the forms of an e-process and a p-process requires calibrating one to another. Prior work by Shafer et al., (2011); Vovk and Wang, (2021) suggests that any e-process can be calibrated into a p-process and vice versa. Hereafter, we allow e-processes and calibrators to be defined on the extended nonnegative reals, $0,\infty]$, and we refer the reader to Wang and Ramdas, 2025b (, section 2.2) for the arithmetic and measure-theoretic properties on $0,\infty]$.
@@
In the opposite direction, p-to-e calibrators (Vovk,, 1993; Sellke et al.,, 2001; Shafer et al.,, 2011; Ramdas et al.,, 2020; Vovk and Wang,, 2021) allow us to convert a p-process into an e-process. A decreasing function $\mathsf{C}:[0,1]\to[0,\infty]$ is a p-to-e calibrator if
@@
and it is admissible (i.e., no other p-to-e calibrator strictly dominates it) if and only if it is right-continuous, $\mathsf{C}(0)=\infty$, and the condition (9) holds with equality (Vovk and Wang,, 2021, ppn. 2.1). Since there is no benefit to using a non-admissible p-to-e calibrator, we restrict our attention to admissible ones hereafter, unless specified otherwise. As its name suggests, a p-to-e calibrator $\mathsf{C}$ converts any p-value $p$ for some null $P$ into an e-value $\mathsf{C}(p)$ for $P$, and analogously, any p-process $(\mathfrak{p}_{t})_{t\geq 0}$ for $\mathcal{P}$ into an e-process $(\mathsf{C}(\mathfrak{p}_{t}))_{t\geq 0}$ for $\mathcal{P}$ (Ramdas et al.,, 2020, ppn. 14).
@@ 2.3 Adjusting the running maximum of e-processes
In general, the running maximum of a test martingale or an e-process is not an e-process (Ramdas et al.,, 2022); in fact, the running maximum of a test martingale, in expectation under the null, can tend to infinity (we describe a concrete example in Section D). For test (super)martingales, Dawid et al., 2011a ; Dawid et al., 2011b ; Shafer et al., (2011) develop a class of functions called adjusters, which convert the running maximum of a test (super)martingale into an e-process. The papers use the terms simple lookback adjusters, capital calibrators, and martingale calibrators, respectively; here, we call them adjusters for brevity. In the context of statistical inference, adjusters allow “insuring against the loss of evidence” by trading off some amount of evidence. This is useful when, e.g., the test martingale sharply decreases following a changepoint.
@@
One known class of admissible adjusters is the family of functions $\mathsf{A}_{\kappa}(\mathfrak{e})=\kappa\mathfrak{e}^{1-\kappa}$, for $\kappa\in(0,1)$ (Shafer et al.,, 2011). Following the analogous strategy for p-to-e calibrators, we can derive a hyperparameter-free admissible adjuster by mixing over $\kappa$: $\mathsf{A}_{\mathsf{mix}}(1)=1/2$ and
@@
Other examples include $\mathsf{A}_{\mathsf{KV}}(\mathfrak{e})=\frac{\mathfrak{e}^{2}\log 2}{(1+\mathfrak{e})\log^{2}(1+\mathfrak{e})}$ (Koolen and Vovk,, 2014) and $\mathsf{A}_{\mathsf{sqrt}}(\mathfrak{e})=\sqrt{\mathfrak{e}}-1$.
In Figure 2, we plot four admissible adjusters (in logarithmic scale): $\mathsf{A}_{\mathsf{mix}}$, $\mathsf{A}_{\mathsf{KV}}$, $\mathsf{A}_{\mathsf{sqrt}}$, and $\mathsf{A}_{\kappa}$ for $\kappa=1/2$. While all of these adjusters are admissible, it is apparent some grow substantially faster than others: $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ grow linearly in $\mathfrak{e}$, up to logarithmic factors, and thus they eventually dominate the other three variants, which only grow sublinearly in $\mathfrak{e}$. $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ are close to each other, although $\mathsf{A}_{\mathsf{mix}}$ dominates $\mathsf{A}_{\mathsf{KV}}$ for $\mathfrak{e}\geq 4.40$. Thus, we use the mixture adjuster $\mathsf{A}_{\mathsf{mix}}$ as a default. In Section E.1, we discuss another variant due to Shafer et al., (2011) that grows faster than $\mathsf{A}_{\mathsf{mix}}$ by a constant factor by completely zeroing out small values of $\mathfrak{e}$.
@@
Note that, even when we use $\mathsf{A}_{\mathsf{mix}}$, we pay a logarithmic “price” relative to the original e-process. Nevertheless, for any admissible adjuster $\mathsf{A}$, we must have $\lim_{\mathfrak{e}\to\infty}\mathsf{A}(\mathfrak{e})=\infty$, implying that the adjusted version of a powerful e-process is still powerful.
@@ 3.1 Combining with a p-process in a sub-filtration via $\mathfrak{p}$-lifting
We first establish that any p-process can be lifted [^2] This notion of lifting across filtrations is largely distinct from the notion of lifting in early studies of complex-valued, continuous-time martingales (Getoor and Sharpe,, 1972). See Section F for additional context. into a finer filtration without calibration. This somewhat surprising result is a direct consequence of the “equivalence lemma” (Ramdas et al.,, 2020; Howard et al.,, 2021) between time-uniform validity, random time validity, and anytime-validity of sequential tests. Specifically, given a p-process $(\mathfrak{p}_{t})_{t\geq 0}$ for a null hypothesis $\mathcal{P}$ in a filtration $\mathbb{F}$, the lemma says that the following statements involving the probability of erroneously rejecting the null are equivalent, given any $\alpha\in(0,1)$ and for each $P\in\mathcal{P}$:
@@
    For any $\mathbb{G}$-stopping time $\tau$, $P(\xi_{\tau})\leq\alpha$.
@@
    For any $\mathbb{F}$-stopping time $\tau$, $P(\xi_{\tau})\leq\alpha$.
@@
Thus, $\mathfrak{p}$-lifting is “free”: any p-process can be lifted without calibration. Theorem 1 follows by applying Lemma 1 to the sequence of $\mathbb{G}$-adapted events, $\xi_{t}=\{\mathfrak{p}_{t}\leq\alpha\}$ for each $t$.
We note in passing that, by complete analogy to Theorem 1, we can also “freely” lift sequential tests and even confidence sequences (CS; Darling and Robbins,, 1967; Howard et al.,, 2021) at arbitrary stopping times. However, we do not focus on them in this work.
The p-lifting result (Theorem 1) tells us how we can combine an e-process $\mathfrak{e}$ in $\mathbb{F}$ with a p-process $\mathfrak{p}$ from a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$. We can first lift the p-process $\mathfrak{p}$ into the $\mathbb{F}$; calibrate it into an e-process $\mathsf{C}(\mathfrak{p})$ in $\mathbb{F}$ via a p-to-e calibrator $\mathsf{C}$; and then combine it with the e-process $\mathfrak{e}$, using the fact that any set of e-processes in the same filtration can be combined into an e-process via averaging (Vovk and Wang,, 2021). This is summarized in the following corollary.
@@ 3.2 Combining with an e-process in a sub-filtration via $\mathfrak{e}$-lifting
Next, we consider the problem of combining two e-processes, one residing in a sub-filtration of another’s filtration. Recall Example 1, where we saw that simply taking the average does not yield an e-process in the finer filtration. As we remarked in Section 3.1, we now know this is because e-processes, whose guarantee relies on an expectation, do not enjoy the same “free” lifting property as p-processes, whose guarantee relies on a probability.
@@
Theorem 2 establishes that, if $\mathfrak{e}$ is anytime-valid in a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$, then $\mathfrak{e}^{\mathsf{adj}}$ is anytime-valid in the finer filtration $\mathbb{F}$. This implies that we can always lift the anytime-validity of an e-process to the finer filtration via adjusters. The theorem also says an adjuster can lift an e-process to any finer filtration $\mathbb{F}$ since the choice of $\mathbb{F}$ was arbitrary. In particular, if an e-process is constructed in a coarsening of the data filtration, then the adjusted process is anytime-valid in the data filtration, so it is valid at any data-dependent sample sizes. The adjusted process is even valid in any enlargement of the data filtration ($\mathbb{F}^{\prime}\supseteq\mathbb{F}$), such that if the e-process we want to combine it with involves other (independent) random variables, say for external randomization, then the adjusted e-process is valid at stopping times involving them.

### v5

@@ 1.3 Related work on combining e-values (and p-values)
Many recent papers highlight how e-values, unlike p-values, can be combined seamlessly, even when they are arbitrarily dependent (e.g., Vovk and Wang, 2021; Wang and Ramdas, 2022; Grünwald et al., 2024). In particular, Vovk and Wang (2021) show that the arithmetic mean function, which straightforwardly yields a valid combined e-value, dominates any other symmetric “e-merging” function. This is useful when combining two or more different e-values for a hypothesis $\mathcal{P}$ and each e-value is constructed using the same data sample $X_{1},\dotsc,X_{n}$. In the sequential context, given that an e-process in a filtration $\mathbb{F}$ is an e-value at any $\mathbb{F}$-stopping time, any two e-processes in the same filtration $\mathbb{F}$ for a (shared) null hypothesis $\mathcal{P}$ can also be combined seamlessly by averaging. The main challenge we highlight in this work is when two e-processes are defined in two different filtrations, such that one (or both) e-processes must first be lifted into a common filtration before the averaging step.
Much of the literature also concerns cases where the e-values exhibit various dependence structures that can be exploited. Generally, taking the product of e-values is a (weakly) dominating e-merging strategy (Vovk and Wang, 2021, proposition 4.2) when the e-values either are independent or form a supermartingale. Taking the product is a useful strategy in sequential settings involving continuous monitoring (Howard et al., 2021; Waudby-Smith and Ramdas, 2024) and in meta-analysis (ter Schure and Grünwald, 2022; Grünwald et al., 2024). In this work, we focus on cases where the two e-processes are constructed using the same data sequence, so the product strategy does not apply. Others study e-values for multiple hypotheses with different, including arbitrary, dependence structures (e.g., Xu et al., 2021; Wang and Ramdas, 2022; Ignatiadis et al., 2024). We do not generally consider such multiple testing scenarios.
Finally, there are various strategies for combining p-values (see, e.g., Vovk and Wang, 2020). Yet, Vovk et al. (2022)’s recent results suggest that any admissible combination function for p-values must go through e-values. Thus, we primarily focus on e-processes, which are also the fundamental notion of evidence in SAVI (Ramdas et al., 2020, 2023). That said, we will additionally discuss the combination of e-processes with p-processes, which are the anytime-valid counterpart to p-values, as a preliminary step to our main result; this task is related to Ignatiadis et al. (2024)’s study on combining (non-sequential) e-values and p-values for multiple testing. As we shall see later, p-processes can be lifted “freely” across filtrations, and they appear in an intermediary step when combining e-processes across filtrations. The underlying connection is made via p-to-e calibrators, which we review in Section 2.2.
@@ 2.1 SAVI part 2: P-processes, sequential tests, and Ville’s inequality
Continuing with our review of SAVI from Section 1.1, we now define p-processes, which are the sequential generalization of p-values. Given a family of distributions $\mathcal{P}$ and a filtration $\mathbb{F}$, we say that a $[0,1]$-valued $\mathbb{F}$-process $(\mathfrak{p}_{t})_{t\geq 0}$ is a p-process for $\mathcal{P}$ in $\mathbb{F}$ if, for any $\alpha\in(0,1)$,
@@
A p-process readily yields the level-$\alpha$ sequential test $\phi_{t}=\mathds{1}\left(\mathfrak{p}_{t}\leq\alpha\right)$ for each $\alpha\in(0,1)$. An e-process also yields a sequential test via Ville’s inequality (Ville, 1939), which was originally shown for test supermartingales but has since been extended to e-processes (Ramdas et al., 2020). The inequality states that, for any e-process $\mathfrak{e}=(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in $\mathbb{F}$, the probability under $\mathcal{P}$ that $\mathfrak{e}$ exceeds a large threshold is small at any $\mathbb{F}$-stopping time: for any $\alpha\in(0,1)$,
@@
the same filtration (Vovk and Wang, 2020)
@@
Table 1: A comparison of evidence measures in SAVI: e-processes and p-processes. Both the e-process $(\mathfrak{e}_{t})_{t\geq 0}$ and the p-process $(\mathfrak{p}_{t})_{t\geq 0}$ are adapted to some underlying filtration $\mathbb{F}$, in which a stopping time $\tau$ can be defined. In the first row, $P\in\mathcal{H}_{0}$ denotes any distribution in the composite null hypothesis $\mathcal{H}_{0}$. The last row summarizes the main takeaways of Section 3.
@@ 2.2 Calibrating e-processes and p-processes
Combining evidence in the forms of an e-process and a p-process requires calibrating one to another. Prior work by Shafer et al. (2011); Vovk and Wang (2021) suggests that any e-process can be calibrated into a p-process and vice versa. Hereafter, we allow e-processes and calibrators to be defined on the extended nonnegative reals, $0,\infty]$, and we refer the reader to Wang and Ramdas (2025b, section 2.2) for the arithmetic and measure-theoretic properties on $0,\infty]$.
@@
In the opposite direction, p-to-e calibrators (Vovk, 1993; Sellke et al., 2001; Shafer et al., 2011; Ramdas et al., 2020; Vovk and Wang, 2021) allow us to convert a p-process into an e-process. A decreasing function $\mathsf{C}:[0,1]\to[0,\infty]$ is a p-to-e calibrator if
@@
and it is admissible (i.e., no other p-to-e calibrator strictly dominates it) if and only if it is right-continuous, $\mathsf{C}(0)=\infty$, and the condition (9) holds with equality (Vovk and Wang, 2021, ppn. 2.1). Since there is no benefit to using a non-admissible p-to-e calibrator, we restrict our attention to admissible ones hereafter, unless specified otherwise. As its name suggests, a p-to-e calibrator $\mathsf{C}$ converts any p-value $p$ for some null $P$ into an e-value $\mathsf{C}(p)$ for $P$, and analogously, any p-process $(\mathfrak{p}_{t})_{t\geq 0}$ for $\mathcal{P}$ into an e-process $(\mathsf{C}(\mathfrak{p}_{t}))_{t\geq 0}$ for $\mathcal{P}$ (Ramdas et al., 2020, ppn. 14).
@@ 2.3 Adjusting the running maximum of e-processes
Generally, the running maximum of a test martingale or an e-process is not an e-process (Ramdas et al., 2022); in fact, the running maximum of a test martingale, in expectation under the null, can tend to infinity (we describe a concrete example in Section D). For test (super)martingales, Dawid et al. (2011a, b); Shafer et al. (2011) develop a class of functions called adjusters, which convert the running maximum of a test (super)martingale into an e-process. The papers use the terms simple lookback adjusters, capital calibrators, and martingale calibrators, respectively; here, we call them adjusters for brevity. In the context of statistical inference, adjusters allow “insuring against the loss of evidence” by trading off some amount of evidence. This is useful when, e.g., the test martingale sharply decreases following a changepoint.
@@
One known class of admissible adjusters is the family of functions $\mathsf{A}_{\kappa}(\mathfrak{e})=\kappa\mathfrak{e}^{1-\kappa}$, for $\kappa\in(0,1)$ (Shafer et al., 2011). Following the analogous strategy for p-to-e calibrators, we can derive a hyperparameter-free admissible adjuster by mixing over $\kappa$: $\mathsf{A}_{\mathsf{mix}}(1)=1/2$ and
@@
Other examples include $\mathsf{A}_{\mathsf{KV}}(\mathfrak{e})=\frac{\mathfrak{e}^{2}\log 2}{(1+\mathfrak{e})\log^{2}(1+\mathfrak{e})}$ (Koolen and Vovk, 2014) and $\mathsf{A}_{\mathsf{sqrt}}(\mathfrak{e})=\sqrt{\mathfrak{e}}-1$.
In Figure 2, we plot four admissible adjusters (in logarithmic scale): $\mathsf{A}_{\mathsf{mix}}$, $\mathsf{A}_{\mathsf{KV}}$, $\mathsf{A}_{\mathsf{sqrt}}$, and $\mathsf{A}_{\kappa}$ for $\kappa=1/2$. While all of these adjusters are admissible, it is apparent some grow substantially faster than others: $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ grow linearly in $\mathfrak{e}$, up to logarithmic factors, and thus they eventually dominate the other three variants, which only grow sublinearly in $\mathfrak{e}$. $\mathsf{A}_{\mathsf{mix}}$ and $\mathsf{A}_{\mathsf{KV}}$ are close to each other, although $\mathsf{A}_{\mathsf{mix}}$ dominates $\mathsf{A}_{\mathsf{KV}}$ for $\mathfrak{e}\geq 4.40$. Thus, we use the mixture adjuster $\mathsf{A}_{\mathsf{mix}}$ as a default. In Section E.1, we discuss another variant due to Shafer et al. (2011) that grows faster than $\mathsf{A}_{\mathsf{mix}}$ by a constant factor by completely zeroing out small values of $\mathfrak{e}$.
@@
We remark that, even when we use $\mathsf{A}_{\mathsf{mix}}$, we pay a logarithmic “price” relative to the original e-process. Nevertheless, for any admissible adjuster $\mathsf{A}$, we must have $\lim_{\mathfrak{e}\to\infty}\mathsf{A}(\mathfrak{e})=\infty$, implying that the adjusted version of a powerful e-process is still powerful.
@@ 3.1 Combining with a p-process in a sub-filtration via $\mathfrak{p}$-lifting
We first establish that any p-process can be lifted [^1] This notion of lifting across filtrations is largely distinct from the notion of lifting in early studies of complex-valued, continuous-time martingales (Getoor and Sharpe, 1972). See Section F for additional context. into a finer filtration without calibration. This somewhat surprising result is a direct consequence of the “equivalence lemma” (Ramdas et al., 2020; Howard et al., 2021) between time-uniform validity, random time validity, and anytime-validity of sequential tests. Specifically, given a p-process $(\mathfrak{p}_{t})_{t\geq 0}$ for a null hypothesis $\mathcal{P}$ in a filtration $\mathbb{F}$, the lemma says that the following statements involving the probability of erroneously rejecting the null are equivalent, given any $\alpha\in(0,1)$ and for each $P\in\mathcal{P}$:
@@
    For any arbitrary $\mathbb{G}$-stopping time $\tau$, $P(\xi_{\tau})\leq\alpha$.
@@
    For any arbitrary $\mathbb{F}$-stopping time $\tau$, $P(\xi_{\tau})\leq\alpha$.
@@
That is, $\mathfrak{p}$-lifting is “free”: any p-process can be lifted without calibration. Theorem 1 follows by applying Lemma 1 to the sequence of $\mathbb{G}$-adapted events, $\xi_{t}=\{\mathfrak{p}_{t}\leq\alpha\}$ for each $t$.
We note in passing that, by complete analogy to Theorem 1, we can also “freely” lift sequential tests and even confidence sequences (CS; Darling and Robbins, 1967; Howard et al., 2021) at arbitrary stopping times. However, we do not focus on them in this work.
The p-lifting result (Theorem 1) tells us how we can combine an e-process $\mathfrak{e}$ in $\mathbb{F}$ with a p-process $\mathfrak{p}$ from a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$. We can first lift the p-process $\mathfrak{p}$ into the $\mathbb{F}$; calibrate it into an e-process $\mathsf{C}(\mathfrak{p})$ in $\mathbb{F}$ via a p-to-e calibrator $\mathsf{C}$; and then combine it with the e-process $\mathfrak{e}$, using the fact that e-processes in the same filtration can be combined into an e-process via averaging (Vovk and Wang, 2021). This is summarized in the following corollary.
@@ 3.2 Combining with an e-process in a sub-filtration via $\mathfrak{e}$-lifting
Next, we consider the problem of combining two e-processes, one residing in a sub-filtration of another’s filtration. Recall Example 1, where we saw that simply taking the average does not yield an e-process in the finer filtration. As we remarked in Section 3.1, we now know this is because e-processes, whose guarantee relies on an expectation bound, do not enjoy the same “free” lifting property as p-processes, whose guarantee relies on a probability bound.
@@
Theorem 2 establishes that, if $\mathfrak{e}$ is anytime-valid in a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$, then $\mathfrak{e}^{\mathsf{adj}}$ is anytime-valid in the finer filtration $\mathbb{F}$. This implies that we can always lift the anytime-validity of an e-process to the finer filtration via adjusters. The theorem also says an adjuster can lift an e-process to any finer filtration $\mathbb{F}$ since the choice of $\mathbb{F}$ was arbitrary for any given $\mathbb{G}$. In particular, if an e-process is constructed in a coarsening of the data filtration, then the adjusted process is anytime-valid in the data filtration, so it is valid at any data-dependent sample sizes. The adjusted process is even valid in any enlargement of the data filtration ($\mathbb{F}^{\prime}\supseteq\mathbb{F}$), such that if the e-process we want to combine it with involves other (independent) random variables, say for external randomization, then the adjusted e-process is valid at stopping times involving them.